OpenStudy (anonymous):

To use a vacuum cleaner at a car wash, customers must pay an initial $3.00 and $0.25 for each 15 min interval. Which table shows how much it costs to use the vacuum cleaner for different amounts of time?

OpenStudy (anonymous):

A. Vacuum Cleaner Costs 15 min intervals 3 5 7 9 Cost $0.75 $1.25 $1.75 $2.25 B. Vacuum Cleaner Costs 15 min intervals 3 5 7 9 Cost $3.75 $4.25 $4.75 $5.25 C. Vacuum Cleaner Costs 15 min intervals 3 5 7 9 Cost $14.25 $20.50 $29.25 $36.75 D. Vacuum Cleaner Costs 15 min intervals 3 5 7 9 Cost $15.75 $16.25 $16.75 $17.25

OpenStudy (anonymous):

@MarrisaAthey

OpenStudy (anonymous):

I was think A or B but then the $3.00 only comes once so i'm not very sure..... but if it was me I would do A

OpenStudy (anonymous):

Can you help me with a few more?

OpenStudy (anonymous):

Sure.. I will try/do my best

OpenStudy (anonymous):

Which situation does the graph represent? A. Holly walks 0.05 mi/min and will complete a 6 mi walk. The graph shows the remaining distance at different times during the walk. B. Holly walks 4 mi/min and will complete a 120 mi walk. The graph shows the remaining distance at different times during the walk. C. Holly walks 6 mi/min and will complete a 20 mi walk. The graph shows the remaining distance at different times during the walk. D. Holly walks 40 mi/min and will complete a 120 mi walk. The graph shows the remaining distance at different times during the walk.

OpenStudy (anonymous):

I believe it would be B

OpenStudy (anonymous):

Taryn conducted a science experiment on saturation. She added sugar to a sugar-water solution at different intervals. The graph shows how much sugar was in the solution at different times. Complete the equation that represents Taryn's situation. y=_______x+_________

OpenStudy (anonymous):

y= 0.5x+1 I think I really wasn't that sure on this one

OpenStudy (anonymous):

Two more

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Eric has a gift card to his favorite restaurant. Every Tuesday he has the same meal at the restaurant. What is the equation that represents Eric's gift card balance? (y - 100) =______(x -______) Eric's Gift Card Weeks 2 4 6 8 Balance $125 $100 $75 $50

OpenStudy (anonymous):

(y - 100) = 25(x - 4) I belive

OpenStudy (anonymous):

Last one!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@Maddox1000